geerky42 (geerky42):

Does Batman Equation really work? And can we "union" relations? (Sorry for long reply.)

geerky42 (geerky42):

I just happened to discover this question in Math Stack Exchange (MSE): http://math.stackexchange.com/questions/54506/is-this-batman-equation-for-real/54568#54568 Here's the image:  http://i.stack.imgur.com/VYKfg.jpg I tried to graph it on Desmos: https://www.desmos.com/calculator/cscx2zcrlf Graphing each factor separately turns out well, however when I multiply all factors, graph won't appear. I think Desmos just choke on it, so I tried something simple; https://www.desmos.com/calculator/enxuzekis6 Yet it doesn't show anything... So apparently image above is not true. My questions are: \(\Large 1)\) Is there any way to union relations? By "union," I mean in \(\mathbb R^2\), if you union \(f(x,y)=0\) and \(g(x,y)=0\), then you would have new union relation \(U(x,y)=0\), which if you graph it, then \(f(x,y)=0\) and \(g(x,y)=0\) would both be graphed simultaneously. From image, it's \(U(x,y) = f(x,y)g(x,y)=0\), but it seems that it is not true. \(\Large 2)\) Why does no one seem to notice that image is not true in MSE question?

OpenStudy (anonymous):

A relation is a set, so obviously you could union two relations.

geerky42 (geerky42):

Yeah I know. Not sure what's correct word for that, but let's just stick with "union" By "union," I mean saying if you graph \(f(x,y)=0\) and \(g(x,y)=0\), you would see something on xy plane. If we "union" these relations, then we would have new relation \(U(x,y)=0\), such that if you graph \(U(x,y)\), then you would see same thing as if you graph \(f(x,y)=0\) and \(g(x,y)=0\) separately at the same time.

geerky42 (geerky42):

From shown image, you just multiply relations, but to me, it doesn't work.

OpenStudy (just_one_last_goodbye):

In my opinion it doesn't work :/ very complex and in testing ur not going to really make a perfect batman sign on the graph xD

OpenStudy (anonymous):

We are such math nerds XD

geerky42 (geerky42):

For example; saying we have \(f(x,y) = x^2+y^2 -1= 0\) and \(g(x,y) = (x+2)^2+y^2-1=0\) Then \(U(x,y)\) would graph this: Exactly what is \(U(x,y)\)? and how can I "combine" f(x,y) and g(x,y) to form U(x,y)?